ROTATIONAL PREDISSOCIATION IN THE GROUND X^{2}Σ^{+} STATE OF ZnH

\underline{T. HIRAO}, K. TERESZCHUK, and P. F. BERNATH, Department of Chemistry, University of Waterloo, Waterloo, Ontario, CANADA N2L 3G1.

It is known that ZnH has a relatively shallow potential curve in its ground state (D_{e} ∼ 7, 660 cm^{−1}), however, accurate spectral information on the higher vibrational levels in the X^{2}Σ^{+} state is not available. In this study, we recorded Fourier transform spectra in 400-500 nm region using a DC discharge inside a tube furnace. A deep blue emission was observed over the granulated zinc heated to 770 K in a slow flow of Ar (∼ 5 Torr) with a trace of H_{2}. So far we identified A^{2}Π − X^{2}Σ^{+} transitions up to v′ = 2 and v′′ = 4, as well as the B^{2}Σ^{+} − X^{2}Σ^{+} 0-4 band. In this B − X band, lines are found to be diffuse around N′′ = 21 and completely dissapear for higher-N′′. Based on the potential curve in the ground state, we found that this happens because of rotational predissociation. We also discuss a deperturbation of the A^{2}Π and B^{2}Σ^{+} states.